A current of one ampere is passed through a straight wire of length 2 metre. The magnetic field at a point in air at a distance of 3 m from one end of the wire but lying on the axis of the wire will be



1
Expert's answer
2021-08-31T17:12:35-0400

By Biot-Savart Law

"dB=\frac{\mu_o I}{4\pi}(\overrightarrow {dl} \times \overrightarrow{r})"

"\overrightarrow {dl} \times \overrightarrow{r}=0" as they are collinear.

Hence, the magnetic field is zero